Find the general solutions of the following differential equations using D-operator methods:

(D^2 - 2D + 5) y = x + 5



1
Expert's answer
2022-03-14T18:34:36-0400

The auxiliary equation is, "m^2 -2m+5 = 0".


Solving for m, "m = \dfrac{2\pm\sqrt{4 - 4\cdot 1 \cdot 5}}{2} = \dfrac{2\pm\sqrt{-16}}{2} = 1\pm2i"


The complementary function is, "\text{C.F} = e^x(c_1 \cos 2x + c_2 \sin 2x)"


"\begin{aligned}
\text{Particular Integral} &= \frac{1}{D^2-2D+5} (x+5)\\
&=\dfrac{1}{5(1+(\frac{D^2 - 2D}{5}))}(x+5) \\
&=\dfrac{1}{5}\left(1+(\frac{D^2 - 2D}{5})\right)^{-1} (x+5)\\
&=\dfrac{1}{5}\left(1 - (\frac{D^2 - 2D}{5}) + (\frac{D^2 - 2D}{5})^2 +\cdots \right)(x+5)\\
&=\dfrac{1}{5}\left(1 - (\frac{D^2 - 2D}{5})\right)(x+5)~~~(\text{Neglecting higher differentials})\\
&=\dfrac{1}{5}\Big((x+5) - \frac{1}{5}\left(D^2(x+5) - 2D(x+5)\right)\Big)\\
&=\dfrac{1}{5}\Big(x+5 + \frac{2}{5}\Big)\\
\therefore P.I &=\dfrac{1}{25}\Big(5x+ 27\Big)\\ 
\end{aligned}"


The general solution is, "y= e^x(c_1 \cos2x + c_2 \sin 2x)+ \dfrac{5x+27}{25}"